# Turning Trash into Treasure: Accelerating Inference of Large Language Models with Token Recycling

### v1

@@ 5.2 Ablation Study
Figure 4: Time allocation for each operation when LLMs response to a query.
5.3 Time Allocation
@@
6 Related Work
Efficient inference is crucial for real-time applications and low-resource scenarios. Many strategies have been developed to reduce latency (Zhou et al. 2024b). Among these, speculative decoding (Chen et al. 2023; Leviathan, Kalman, and Matias 2023; Miao et al. 2024; Xia et al. 2023) has emerged as a losses technique for achieving lossless speedup by predicting multiple possible continuations simultaneously, thus reducing the number of decoding steps needed without compromising accuracy. Some speculative decoding methods rely on additional draft models to guess draft tokens. These typically involve using smaller models from the same series (Zhao et al. 2024; Spector and Re ; Sun et al. 2023; Liu et al. 2024b; Yuan et al. 2024; Gong et al. 2024) or training new models with a shared vocabulary (Leviathan, Kalman, and Matias 2023; Chen et al. 2023; Zhou et al. 2024a; Li et al. 2024). Additionally, Kou et al. (2024); Wang et al. (2024b) propose training the original LLMs to enable non-aggressive decoding. While effective, these approaches require managing or training multiple models, which can be non-trivial and resource-intensive. Other methods focus on parameter-efficient structures. These approaches minimize the need for complete retraining but still require model-specific training and adaptation, limiting their scalability and general applicability (Lin et al. 2024; Liu et al. 2024a).
There are also train-free methods that construct retrieval libraries to obtain draft tokens (Yang et al. 2023). Lookahead (Fu et al. 2024) uses the LLM to generate n-grams through multiple decodings, building a retrieval library that can hit multiple tokens in one step. Despite this advantage, it requires the LLM to generate n-grams while responding to queries, which reduces efficiency. PLD (Saxena 2023) retrieves only from previous content, resulting in minimal overhead and significant speedup in high-redundancy tasks like summarization. However, it provides little acceleration for tasks requiring the generation of new content, like translation. REST (He et al. 2023) constructs retrieval libraries using existing corpora and performs well in common scenarios. However, this approach requires large storage, time-consuming retrieval, and cannot adapt to each query.
Token Recycling is a train-free, retrieval-based method that stores candidate tokens in an adjacency matrix and reuses them as draft tokens. It requires no additional generation, covers a broader range of possible continuations, and demands minimal storage with low retrieval costs. The update process ensures the retrieval space remains adaptable
7 Conclusion
In this work, we introduce Token Recycling, a speculative decoding method for accelerating the inference of LLMs. It utilizes an adjacency matrix to store candidate tokens and retrieve a draft tree, which is then verified with tree attention. The matrix is updated with new candidate tokens generated during decoding. Token Recycling could be integrated seamlessly with existing LLMs and various tasks. As a train-free approach, it achieves a speedup of approximately 2x with <2MB of additional storage, improving over 31% compared to previous train-free approaches.

### v2

@@ A.3 Time Allocation
Figure 4: Time allocation for each operation when LLMs respond to a query.
@@
A.4 Compare With Eagle
TR is Self-Draft and Train-Free, and fair comparisons with other speculative decoding methods that do not require training (e.g., Lookahead, PLD, REST) have already been conducted in section 4.2. However, due to the effectiveness of our approach, our performance significantly exceeds that of these baselines. As a result, we became curious about comparing our method with training-dependent approaches, such as representively, Medusa. The comparison results show that our method even outperforms it a lot. Approaches like Eagle Li et al. (2024), which rely on additional architectures to generate drafts, may achieve higher performance, but they require collecting datasets for training the small model, and during deployment, they also need extra memory. In contrast, TR only requires <2MB of storage, is train-free, and can be directly applied to all LLMs, achieving nearly a 2x speedup across various model sizes. In Table 6, we compare the performance of Token Recycling and Eagle on the vicuna-7b-v1.3 model using SpecBench. In comparison, our method uses only 0.39% of the memory required by Eagle, while achieving 88.65% of its speedup performance.
Memory Cost Speed Ratio
Eagle 500MB 2.29
TR <2MB 2.03
Table 6: Compare Token Recycling with Eagle.
A.5 Reuse Mechanism Analysis
The output of LLM is context-dependent, it makes us curious why reusing candidate tokens from previous generations works. Xiao et al. (2024) analyzed the distribution of attention logits in Transformers and found that the first two layers focus more on ‘local’ patterns, with recent tokens receiving much more attention. In later layers, the model shifts its focus to the tokens at the beginning of the sequence. To accelerate, candidate tokens need to satisfy both local semantics and long-range dependencies. As shown in Figure 1, draft tokens are divided into accepted and rejected tokens. Accepted tokens are those that appear in the previous query response. In Token Recycling, candidate tokens for all draft tokens are stored, not only accepted ones. Each decoding step involves 79 draft tokens, meaning 79 tokens receive/update their candidate tokens at each step. The quantity of rejected tokens is far more than accepted tokens, and their candidate tokens are also stored in the matrix. In other words, previous generations actually provide a large number of common patterns stored in the matrix, and these patterns often meet local semantic needs.
From two perspectives, the reuse of these common patterns is justified:
For scenarios like sentence transitions, verb collocations, punctuation, or words split into multiple tokens, there is often no need for long-range dependencies. The common patterns stored in the matrix can significantly accelerate the decoding process.
Rejected tokens may not satisfy the long-range dependencies of the previous, but this does not mean they do not meet the long-range dependencies of the current. These tokens may be accepted in the future. As shown in Table 4, the performance gain from the candidate tokens of rejected tokens is significantly greater than the gain from using only the accepted tokens in the SpecBench experiments. We also include a case study in A.6.
A.6 Case Study
Below, we present a real example from MT-Bench, illustrating how accepted tokens and rejected tokens contribute to the acceleration of Token Recycling. First Round:
  • Prefix last token: [’guest’]
  • Merged sequence with draft tokens: [’guest’, ’speaker’, ’</s>’, ’speak’, ’<0x0A>’, ’Spe’, ’lect’, ’speaking’, ’spe’, ’at’, ’for’, ’</s>’, ’is’, ’could’, ’can’, ’would’, ’<0x0A>’, ’<s>’, ’sime’, ’multicol’, ’bolds’, ’</s>’, ’ings’, ’in’, ’<0x0A>’, ’a’, ’aking’, ’ures’, ’engag’, ’aking’, ’a’, ’an’, ’the’, ’[’, ’our’, ’local’, ’</s>’, ’up’, ’a’, ’an’, ’</s>’, ’<s>’, ’sime’, ’a’, ’be’, ’help’, ’like’, ’<0x0A>’, ’The’, ’Home’, ’guest’, ’<s>’, ’<0x0A>’, ’opportunity’, ’</s>’, ’local’, ’nearby’, ’guest’, ’up’, ’public’, ’guest’, ’</s>’, ’guest’, ’public’, ’local’, ’local’, ’The’, ’first’, ’The’, ’<0x0A>’, ’event’, ’community’, ’Toast’, ’Buddh’, ’speaker’, ’speaker’, ’event’, ’time’, ’.’, ’,’, ’ism’]
  • Accepted sequence: [’guest’, ’speaker’, ’at’, ’a’, ’local’, ’event’, ’could’]
Key observations:
  • Rejected tokens still receive candidate tokens: For example, [’be’] was not directly accepted, but its candidate tokens were stored.
  • Candidate tokens of accepted tokens are stored: [’local’] for [’a’] was chosen in this round, but other candidate tokens were also retained.
At this point, the adjacency matrix stores:
  • ’be’: [’able’, ’onto’, ’mistaken’, ’wrong’, ’the’, ’interested’, ’persu’, ’a’]
  • ’a’: [’local’, ’time’, ’personal’, ’professional’, ’few’, ’unique’, ’great’, ’low’]
Second Round:
  • Current prefix last token: [’could’]
  • Merged sequence with draft tokens: [’could’, ’be’, ’provide’, ’</s>’, ’help’, ’offer’, ’present’, ’not’, ’actually’, ’a’, ’an’, ’the’, ’</s>’, ’just’, ’one’, ’benef’, ’both’, ’,’, "’", ’you’, ’for’, ’<s>’, ’sime’, ’multicol’, ’you’, ’overcome’, ’some’, ’</s>’, ’public’, ’<unk>’, ’great’, ’fant’, ’wonderful’, ’valuable’, ’unique’, ’perfect’, ’</s>’, ’ter’, ’up’, ’event’, ’local’, ’up’, ’local’, ’<s>’, ’not’, ’of’, ’ited’, ’models’, ’such’, ’like’, ’my’, ’The’, ’as’, ’comp’, ’<s>’, ’opportunity’, ’guest’, ’</s>’, ’speaker’, ’way’, ’astic’, ’asy’, ’bl’, ’ins’, ’guest’, ’coming’, ’coming’, ’as’, ’first’, ’a’, ’at’, ’is’, ’would’, ’event’, ’<s>’, ’natural’, ’public’, ’a’, ’a’, ’an’, ’could’]
  • Accepted sequence: [’could’, ’be’, ’a’, ’great’]
At this point:
  • Candidate tokens from rejected tokens are matched: ’be’ correctly predicted ’a’.
  • Candidate tokens from accepted tokens are matched: ’a’ successfully predicted ’great’.
Figure 5: The static tree used in Token Recycling.